YoursTruly:

Given a slope of -6/7 and a y-intercept of -10, write the equation of the line in slope intercept form

coacoapuffprincess:

uM

YoursTruly:

Shimpy help I'm begging.

coacoapuffprincess:

Guive me a second

YoursTruly:

K

YoursTruly:

I'm gonna fail doe

coacoapuffprincess:

Y=MX+b

5sos:

Ok. y=-6/7x-10 -6/7 is the slope aka m -10 is the y-intercept aka b All you have to do is plug everything in.

YoursTruly:

Do that for me please? :)

5sos:

I did. XD

YoursTruly:

TY

tshimp0629:

good job everyone

5sos:

Ok. y=-6/7x-10 <


answer -6/7 is the slope aka m <
explanation -10 is the y-intercept aka b <
explanation All you have to do is plug everything in.<---explanation

5sos:

oml bruh Ok. y=-6/7x-10 answer -6/7 is the slope aka m explanation -10 is the y-intercept aka b explanation All you have to do is plug everything in. explanation